OpenStudy (anonymous):

suppose that you draw 2 cards from a standard deck. How many possible draws are there? what is the probability that you drew a pair of 6's? what are the odds against drawing a pair of 6's?

OpenStudy (anonymous):

How many cards are there?

OpenStudy (anonymous):

...52

OpenStudy (anonymous):

52

OpenStudy (anonymous):

When you draw one card, now how many cards are there?

OpenStudy (anonymous):

...51

OpenStudy (anonymous):

yes 51

OpenStudy (anonymous):

the number of cards is the number of possibilities, and you multiply the possibilities

OpenStudy (anonymous):

52*51 = total possibilities

OpenStudy (anonymous):

so there are 2652 possibilites right...and how can i find out the possibility of drawing a pair of 6's?

OpenStudy (anonymous):

how many sixes are there before you draw any?

OpenStudy (anonymous):

that's what i dont know it does not say

OpenStudy (anonymous):

there are four sixes in a deck

OpenStudy (anonymous):

so 24

OpenStudy (anonymous):

no, the "6" in this case is just a place holder, not a possibility

OpenStudy (anonymous):

so then there would only be 4 possibility because there are only 4 cards of 6's right

OpenStudy (anonymous):

there are 4, then when you draw one, there are 3 to select. So the total possible ways to get your desired outcome is 4*3 = 12 to make this into a probability, divide by the total possibilities (52*51 from before)

OpenStudy (anonymous):

and how would i find the "odd" against drawing a pair of 6's? what does it mean by od

OpenStudy (anonymous):

*odd

OpenStudy (anonymous):

odds means probability. I don't know if odds against is different than odds for something, if it is, then you may have to subtract the answer from 2652/2652 or 1.

OpenStudy (anonymous):

it is either (4*3)/(52*51) = 1/221 or 220/221 if odds against means the odds that you WON'T get a pair of sixes. What do you think it means?

OpenStudy (anonymous):

that wont get a pair of 6's

OpenStudy (anonymous):

anothr quick question: suppose you roll a pair of dice. what is the possibility you roll a total of 9? is it 2 out of 6